# Amir Hossein Fahim Raouf    İsmail Güvenç

### v1

@@ I Introduction
Uncrewed Aircraft Systems (UAS) play an increasingly critical role in mission-critical and commercial applications, including command and control (C2), payload data delivery, and situational awareness. [FotouhiUAVCellularSurvey2019, geraci2022will]. Relative to terrestrial user equipment (UE), UAS operate under fundamentally different propagation conditions, characterized by higher line-of-sight (LoS) probability, increased interference from multiple base stations (BSs), and broader coverage footprints [khuwaja2018survey]. Accurate characterization of altitude-dependent signal behavior is therefore essential for effective interference management, airspace coexistence, and reliable UAS network design [raouf2025wireless].
Most prior work on UAV communications has focused on path loss, shadowing, and fast fading as functions of altitude and horizontal distance, typically within a single-band single-operator context, see e.g., [AlHouraniA2GPathLoss2014, willink2015measurement]. In contrast, practical UAV systems operate in spectrally dense environments where multiple bands and operators coexist and where the spectral structure itself, that is the distribution of power across frequency, varies with altitude. For tasks such as band selection, interference-aware trajectory planning, or RF-based sensing, received power alone is insufficient. It is necessary to characterize how spectral structure and spectrum occupancy evolve with altitude.
Existing spectrum occupancy studies commonly rely on scalar descriptors such as duty cycle, occupancy probability above a fixed threshold, or median and percentile power levels to summarize band usage [yucek2009survey, wellens2007evaluation, LopezBenitezMethodology2010, ChantaveerodFMOccupancy2021, KozlowskiISMOccupancy2021]. While effective for quantifying average utilization, these metrics collapse the internal frequency-domain structure of the received signal and do not capture how power is distributed across subcarriers or how this distribution evolves with altitude. In particular, duty cycle and occupancy probability are highly sensitive to threshold selection and provide no information on spectral dispersion or fragmentation, while median power conflates changes in aggregate power with changes in spectral shape. Consequently, such metrics cannot distinguish between bands with similar average power but fundamentally different spectral organization, nor can they reveal altitude-dependent transitions in interference structure once band-average power has saturated.
@@ II Measurement Campaign and Dataset
As the measurement campaign and associated dataset are described in [maeng2025altitude, raouf2025wireless], only a brief overview of the measurement procedure and dataset is provided here. The payload consisted of a USRP-based software-defined radio (SDR) and a GNSS module providing altitude and time synchronization. The spectrum-monitoring experiment operates by sweeping center frequencies from 87 MHz to 6 GHz in a loop. For each center frequency, the USRP collects 500500k IQ samples at a sampling rate of 30.72 MHz, and Welch’s method is applied using a 512-point FFT. To suppress edge artifacts, FFT bins at the band edges are discarded, and the center frequency is incremented by 25.68 MHz across the sweep. Each sweep is stored for post-processing, and Fig. 1 illustrates the spectrum sweep procedure for the helikite-mounted platform. Although experimental parameters such as USRP gain, sampling rate, number of sweep points, and FFT size are configurable, they were held constant across measurement campaigns to ensure cross-year comparability, with only minor adjustments to experiment duration imposed by operational constraints.
@@ III-A Band-Average Power
This is the conventional band-integrated power, capturing the aggregate measured power from all BSs contributing to band $b$. As altitude increases, $P_{b}(h)$ is expected to exhibit a transition from clutter-limited to LoS-dominated behavior, with a corresponding increase in average received power and decrease in variability [maeng2025altitude].
@@ III-C Spectral Sparsity
Sparsity quantifies the fraction of frequency bins that contain detectable signal energy. Following common practice in spectrum occupancy studies [LopezBenitezMethodology2010, islam2008spectrum], we define a band-specific detection threshold $\gamma_{b}=P_{b,\text{noise}}+\Delta_{\text{th}}$, where $P_{b,\text{noise}}$ is an empirical noise floor estimate based on the fifth percentile of the band’s PSD samples over the entire campaign, and $\Delta_{\text{th}}$ (e.g., 3 dB) is a conservative margin above noise [wellens2007evaluation]. This threshold is fixed per band and campaign year rather than per altitude to avoid confounding sparsity variations with altitude-dependent noise estimation. Similar percentile-based thresholding approaches have been adopted in recent spectrum occupancy studies, for example in FM band measurements [ChantaveerodFMOccupancy2021].
@@ III-D Transition Regions and 10–90% Metrics
  • It robustly captures the effective clutter region associated with rooftops and tree canopies and the altitude range over which LoS probability increases rapidly [AlHouraniA2GPathLoss2014, khuwaja2018survey].
@@ IV-A First-Order Differential Model for Power and Entropy
The first-order structure is motivated by the physical mechanisms governing air-to-ground propagation. In cluttered urban and suburban environments, the most significant changes in received signal characteristics occur as the platform ascends through the altitude range in which the LoS probability increases rapidly. Empirical models show that both LoS probability and excess path loss approach their limiting values approximately exponentially with altitude [AlHouraniA2GPathLoss2014]. Spectral metrics derived from the received signal, including band-average power and spectral entropy, therefore exhibit similar altitude-dependent behavior. Once the platform rises above the effective clutter height, further altitude increases yield diminishing variations, which naturally motivates a first-order relaxation model. The model in (5) is intended to capture the dominant monotonic altitude trends over the measured altitude range and does not account for potential power degradation at very high altitudes, where visibility and path loss effects may reverse the trend.
@@ V-E Limitations and Generalization
It is worth noting that all measurement datasets were calibrated to mitigate receiver-side offsets using free-space path loss analysis of FM broadcast signals. The reported power levels therefore represent calibrated received power at the helikite-mounted SDR and do not correspond to equivalent isotropically radiated power (EIRP), as transmitter-side parameters and propagation losses are unavailable; interested readers are referred to [maeng2025altitude] for further details.
VI Conclusion
@@
This paper introduced ADSSM, a measurement-based framework for characterizing how band-average power, spectral entropy, and sparsity evolve with altitude for signals observed by UAV platforms. The model combines first-order differential equations for power and entropy with a logistic activation model for sparsity, and is physically grounded in the transition from clutter-limited propagation at low altitudes to LoS-dominated conditions as altitude increases. Using multi-year spectrum measurements collected by a tethered UAV across six sub-6 GHz licensed and unlicensed U.S. spectrum allocations, we demonstrated that the ADSSM accurately fits altitude-binned spectral metrics, achieving low root-mean-square error and high coefficients of determination while preserving physically consistent asymptotic behavior. The resulting parameters provide compact and interpretable descriptors of effective clutter height, asymptotic received power, entropy stabilization, and sparsity transition altitudes. The analysis highlights that these transition parameters are environment dependent and should be expected to vary with deployment geometry, terrain, and infrastructure density. Beyond fitting accuracy, the results show that altitude-dependent spectrum behavior is inherently multi-dimensional. In several bands, dominant power transitions occur over substantially narrower altitude ranges than those associated with entropy and sparsity, indicating that interference structure and spectrum occupancy may continue to evolve even after average received power has saturated. This exposes a fundamental limitation of power-only or threshold-based spectrum models commonly used in aerial spectrum sensing. From a dynamic spectrum access perspective, these findings demonstrate that spectrum-aware UAV operation requires explicit modeling of spectral structure, not just received power, particularly in shared and heterogeneous bands where coexistence decisions are sensitive to altitude.

### v2

@@ I Introduction
Uncrewed Aircraft Systems (UAS) play an increasingly critical role in mission-critical and commercial applications, including command and control (C2), payload data delivery, and situational awareness. [3, 4]. Relative to terrestrial user equipment (UE), UAS operate under fundamentally different propagation conditions, characterized by higher line-of-sight (LoS) probability, increased interference from multiple base stations (BSs), and broader coverage footprints [6]. Accurate characterization of altitude-dependent signal behavior is therefore essential for effective interference management, airspace coexistence, and reliable UAS network design [10].
Most prior work on UAV communications has focused on path loss, shadowing, and fast fading as functions of altitude and horizontal distance, typically within a single-band single-operator context, see e.g., [1, 13]. In contrast, practical UAV systems operate in spectrally dense environments where multiple bands and operators coexist and where the spectral structure itself, that is the distribution of power across frequency, varies with altitude. For tasks such as band selection, interference-aware trajectory planning, or RF-based sensing, received power alone is insufficient. It is necessary to characterize how spectral structure and spectrum occupancy evolve with altitude.
Existing spectrum occupancy studies commonly rely on scalar descriptors such as duty cycle, occupancy probability above a fixed threshold, or median and percentile power levels to summarize band usage [14, 12, 8, 2, 7]. While effective for quantifying average utilization, these metrics collapse the internal frequency-domain structure of the received signal and do not capture how power is distributed across subcarriers or how this distribution evolves with altitude. In particular, duty cycle and occupancy probability are highly sensitive to threshold selection and provide no information on spectral dispersion or fragmentation, while median power conflates changes in aggregate power with changes in spectral shape. Consequently, such metrics cannot distinguish between bands with similar average power but fundamentally different spectral organization, nor can they reveal altitude-dependent transitions in interference structure once band-average power has saturated.
@@ II Measurement Campaign and Dataset
As the measurement campaign and associated dataset are described in [9, 10], only a brief overview of the measurement procedure and dataset is provided here. The payload consisted of a USRP-based software-defined radio (SDR) and a GNSS module providing altitude and time synchronization. The spectrum-monitoring experiment operates by sweeping center frequencies from 87 MHz to 6 GHz in a loop. For each center frequency, the USRP collects 500500k IQ samples at a sampling rate of 30.72 MHz, and Welch’s method is applied using a 512-point FFT. To suppress edge artifacts, FFT bins at the band edges are discarded, and the center frequency is incremented by 25.68 MHz across the sweep. Each sweep is stored for post-processing, and Fig. 1 illustrates the spectrum sweep procedure for the helikite-mounted platform. Although experimental parameters such as USRP gain, sampling rate, number of sweep points, and FFT size are configurable, they were held constant across measurement campaigns to ensure cross-year comparability, with only minor adjustments to experiment duration imposed by operational constraints.
@@ III-A Band-Average Power
This is the conventional band-integrated power, capturing the aggregate measured power from all BSs contributing to band $b$. As altitude increases, $P_{b}(h)$ is expected to exhibit a transition from clutter-limited to LoS-dominated behavior, with a corresponding increase in average received power and decrease in variability [9].
@@ III-C Spectral Sparsity
Sparsity quantifies the fraction of frequency bins that contain detectable signal energy. Following common practice in spectrum occupancy studies [8, 5], we define a band-specific detection threshold $\gamma_{b}=P_{b,\text{noise}}+\Delta_{\text{th}}$, where $P_{b,\text{noise}}$ is an empirical noise floor estimate based on the fifth percentile of the band’s PSD samples over the entire campaign, and $\Delta_{\text{th}}$ (e.g., 3 dB) is a conservative margin above noise [12]. This threshold is fixed per band and campaign year rather than per altitude to avoid confounding sparsity variations with altitude-dependent noise estimation. Similar percentile-based thresholding approaches have been adopted in recent spectrum occupancy studies, for example in FM band measurements [2].
@@ III-D Transition Regions and 10–90% Metrics
  • It robustly captures the effective clutter region associated with rooftops and tree canopies and the altitude range over which LoS probability increases rapidly [1, 6].
@@ IV-A First-Order Differential Model for Power and Entropy
The first-order structure is motivated by the physical mechanisms governing air-to-ground propagation. In cluttered urban and suburban environments, the most significant changes in received signal characteristics occur as the platform ascends through the altitude range in which the LoS probability increases rapidly. Empirical models show that both LoS probability and excess path loss approach their limiting values approximately exponentially with altitude [1]. Spectral metrics derived from the received signal, including band-average power and spectral entropy, therefore exhibit similar altitude-dependent behavior. Once the platform rises above the effective clutter height, further altitude increases yield diminishing variations, which naturally motivates a first-order relaxation model. The model in (5) is intended to capture the dominant monotonic altitude trends over the measured altitude range and does not account for potential power degradation at very high altitudes, where visibility and path loss effects may reverse the trend.
@@ V-E Limitations and Generalization
It is worth noting that all measurement datasets were calibrated to mitigate receiver-side offsets using free-space path loss analysis of FM broadcast signals. The reported power levels therefore represent calibrated received power at the helikite-mounted SDR and do not correspond to equivalent isotropically radiated power (EIRP), as transmitter-side parameters and propagation losses are unavailable; interested readers are referred to [9] for further details.
VI Conclusion and Discussion
A spectrum band is appropriate for sharing only if a secondary user can reliably identify idle resources, predict interference well enough to avoid harmful impact, and exploit those opportunities in a sustained and repeatable manner. Satisfying these conditions requires a multidimensional characterization of the interference environment. Aggregated received power provides a first-order measure of how severe interference is and therefore whether operation is even viable. However, power alone is insufficient for sharing decisions, since bands with comparable average power can exhibit fundamentally different internal structures, and low average power does not imply usability if interference is fragmented or unstable. As a result, power must be interpreted together with metrics that describe how interference is distributed across frequency and how predictable it is.
Spectral sparsity and spectral entropy supply this missing information. Sparsity directly quantifies resource availability and is the primary indicator of sharing feasibility, as it captures the fraction of frequency bins that are effectively idle. High sparsity implies abundant transmission opportunities and low collision risk, whereas low sparsity fundamentally limits sharing regardless of average power or predictability. Spectral entropy complements sparsity by characterizing interference structure and predictability. Low entropy indicates dominance by a small number of structured emitters that are easier to sense, model, and avoid, while high entropy reflects many comparable contributors and necessitates conservative access strategies. A share-friendly band therefore exhibits the joint condition of high sparsity, low to moderate entropy, and low to moderate aggregated power, indicating both the presence of idle resources and the ability to exploit them efficiently.
This characterization is particularly relevant for UAVs, which are expected to rely heavily on existing spectrum, including licensed cellular bands and unlicensed ISM bands, for command, control, data links, and electronic conspicuity. As highlighted in recent U.S. Department of Transportation guidance on electronic conspicuity, leveraging existing terrestrial communication infrastructure is a key enabler for safe UAV integration at scale [11]. For UAVs, understanding how aggregated interference power evolves with altitude is essential for quantifying the interference environment experienced in flight, since increasing altitude fundamentally alters propagation and LoS conditions relative to ground users. At the same time, sparsity and entropy provide critical insight into how much spectrum is actually available and how reliably it can be accessed. Together, these metrics enable altitude-aware scheduling and band selection for UAVs, supporting opportunistic yet interference-safe operation that cannot be achieved using power-only or threshold-based models.
@@ VI Conclusion and Discussion
This paper introduced ADSSM, a measurement-based framework for characterizing how band-average power, spectral entropy, and sparsity evolve with altitude for signals observed by UAV platforms. The model combines first-order differential equations for power and entropy with a logistic activation model for sparsity, and is physically grounded in the transition from clutter-limited propagation at low altitudes to LoS-dominated conditions as altitude increases. Using multi-year spectrum measurements collected by a tethered UAV across six sub-6 GHz licensed and unlicensed U.S. spectrum allocations, we demonstrated that the ADSSM accurately fits altitude-binned spectral metrics, achieving low root-mean-square error and high coefficients of determination while preserving physically consistent asymptotic behavior. The resulting parameters provide compact and interpretable descriptors of effective clutter height, asymptotic received power, entropy stabilization, and sparsity transition altitudes. The analysis highlights that these transition parameters are environment dependent and should be expected to vary with deployment geometry, terrain, and infrastructure density. Beyond fitting accuracy, the results reveal that altitude-dependent spectrum behavior is inherently multi-dimensional. In several bands, dominant power transitions occur over narrower altitude ranges than those associated with entropy and sparsity, indicating that interference structure and spectrum occupancy may continue to evolve even after average received power has saturated. This observation underscores the limitations of power-only air-to-ground models for spectrum-aware UAV operation.